OpenStudy (anonymous):

i dont know the slope of 6x+8y+12 can anyone help me ?

OpenStudy (igreen):

I think you wrote your equation wrong. Can you write it again? @poohbear12345

OpenStudy (anonymous):

You can't really get the slope of that question without it being an equation. I would say there is probably a typo and it is supposed to be =12 and not + 12 in that case you just solve for y and and get the coefficient of x/

OpenStudy (anonymous):

it was suppose to be 6x+8y=12

OpenStudy (igreen):

Correction: Slope = \(-\Large\frac{3}{4}\)

OpenStudy (igreen):

You can give medals by clicking 'Best Response'. @poohbear12345

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

whats the y intercept of 3x-5y=6

OpenStudy (anonymous):

To get the y-intercept in a standard form equation you would substitute x for 0 then solve for y, In this case, you would divide 6 by -5 which gives you -5/6